# Empowering Multimodal Respiratory Sound Classification with Counterfactual Adversarial Debiasing for Out-of-Distribution Robustness

### v1

@@ Abstract
Multimodal respiratory sound classification offers promise for early pulmonary disease detection by integrating bioacoustic signals with patient metadata. Nevertheless, current approaches remain vulnerable to spurious correlations from attributes such as age, sex, or acquisition device, which hinder their generalization, especially under distribution shifts across clinical sites. To this end, we propose a counterfactual adversarial debiasing framework. First, we employ a causal graph-based counterfactual debiasing strategy to suppress non-causal dependencies from patient metadata. Second, we introduce adversarial debiasing to learn metadata-insensitive representations and reduce metadata-specific biases. Third, we design counterfactual metadata augmentation to mitigate spurious correlations further and strengthen metadata-invariant representations. By doing so, our method consistently outperforms strong baselines in evaluations under both in-distribution and distribution shifts. The code is available at https://github.com/RSC-Toolkit/BTS-CARD.
@@ 1 Introduction
Respiratory sound classification (RSC) provides a non-invasive and cost-effective method for the early detection of pulmonary diseases, offering a practical alternative to manual auscultation [1, 2, 3]. Clinicians typically base their diagnostic decisions on heterogeneous information sources, including patient metadata [3, 4, 5]. Motivated by this clinical reasoning process, recent advances in RSC have leveraged multimodal learning to integrate respiratory audio with auxiliary metadata (age, sex, etc.) [3].
@@
Causal inference provides a principled framework to address such limitations by estimating the effect of specific variables while controlling for confounders [10, 11]. This perspective is especially valuable in multimodal learning, where multimodal integration can make it more challenging to identify the true causal mechanisms, while also amplifying modality-driven biases. [12, 13]. In RSC, metadata often exhibit skewed distributions across diagnostic categories, leading models to rely on superficial correlations instead of clinically meaningful features [6, 14]. Counterfactual debiasing mitigates this by disentangling causal signals from spurious ones, thereby enhancing in-distribution (ID) performance and out-of-distribution (OOD) robustness [11, 15].
Generalization is critical for clinical AI systems to ensure reliable performance across heterogeneous hospitals and patient populations [14, 16, 17]. Beyond causal approaches, improving robustness under OOD conditions also requires learning representations that remain stable across variations in demographics, recording devices, clinical protocols, etc [18]. Adversarial debiasing, originally introduced for algorithmic fairness, can be adapted to suppress hospital-specific attributes by jointly optimizing the main prediction task and adversarial domain classifiers [19, 20, 21]. In this context, adversarial training facilitates the learning of metadata-agnostic representations, thereby promoting generalization across diverse clinical sites [20]. However, existing research on RSC has overlooked the integration of counterfactual reasoning with adversarial debiasing to improve robustness under distribution shifts.
Therefore, we propose BTS-CARD, a novel counterfactual adversarial debiasing framework for multimodal RSC (see Figure 1). First, we introduce a counterfactual debiasing strategy that suppresses non-causal associations, yielding more reliable predictions in both ID (ICBHI [22]) and OOD (SPRSound [23]) settings. Second, we enhance generalization via adversarial debiasing, which promotes metadata-invariant representations and prevents the model from encoding metadata-specific factors such as acquisition device or recording location. Third, we devise counterfactual metadata augmentation to further mitigate spurious correlations from patient metadata and strengthen metadata-agnostic representations. Extensive experiments demonstrate that our method consistently outperforms strong baselines, validating its robustness and transferability.
Fig. 1: An Overview of Our Counterfactual Adversarial Debiasing Framework for Multimodal RSC.
@@ 2.2 Task Formulation
Debiased Multimodal RSC. From a causal perspective, we decompose the influence of metadata into two pathways: the spurious direct path $t\rightarrow Y_{t,m}$ and the informative indirect path $(a,t)\rightarrow m\rightarrow Y_{t,m}$. Our framework employs counterfactual reasoning, illustrated by contrasting the top (factual) and bottom (counterfactual) graphs in the Counterfactual Inference of Figure 1, to disentangle causal effects from confounding. Combined with adversarial debiasing and counterfactual metadata augmentation, our approach further learns representations resilient to metadata variance, thereby improving predictive accuracy and robustness.
@@ 2.3 Counterfactual Adversarial Debiasing Framework
In this section, we present the counterfactual adversarial framework for multimodal RSC that builds on the BTS model [3].
@@
where $Y_{t,m}$ denotes the prediction under text derived by metadata prompt $t$ [3] and multimodal representation $m$, while $Y_{t^{*},m^{*}}$ corresponds to their counterfactual counterparts. Here, multimodal counterfactual features are instantiated using a dummy vector $m^{*}$, fixed to a constant value of 1. We use RUBi [12] for fusion, given its established effectiveness in causal modeling.
Next, we estimate the NDE of metadata text using Clinical TinyBERT [26], which is not only efficient, but also easily adaptable to new tasks via fine-tuning. The NDE isolates the spurious shortcut $T\to Y$, and the TIE is then obtained by subtracting NDE from TE:
@@
Adversarial Debiasing. To promote metadata-agnostic representations for robustness across different clinical environments, we apply adversarial debiasing [19] to the NDE. We assume that recording device types and locations vary across hospitals and introduce bias during deployment; hence, both are targeted. Debiasing is restricted to the natural direct effect (NDE), with logits $Y_{t}$ fed to an adversary $U$ to yield representation $z$, while classification is optimized with cross-entropy (CE). Thus, with $D_{a}$, the adversarial discriminator for the metadata attribute $a\in\{\text{location},\text{device}\}$ and $l_{a}$, its corresponding ground-truth label, the regarding objective is as:
@@
$\text{where}\quad L_{\text{CE}_{\text{NDE}}}$ $=\text{CE}(Y_{t},y),$
$\text{and}\quad L_{a}$ $=\text{CE}(D_{a}(z),\,l_{a}).$
@@
From a causal perspective, this intervention suppresses the spurious shortcut $T\to Y$ while leaving the TE model intact to preserve both direct and indirect effects [11]. This strategy enforces insensitivity to non-causal attributes without discarding meaningful pathways. It parallels the logic of $do$-calculus [10], where interventions isolate causal effects by severing spurious dependencies. Replacing sensitive metadata in this manner thus mimics a $do$-operation that removes the $T\to Y$ link while retaining valid mediating paths, aligning model behavior with the identifiable causal structure.
@@
Thus, with $\lambda_{\text{CE}}$ and $\lambda_{\text{KL}}$ controlling the weight of CE and KL regularization, the overall training loss is:
@@
At inference, we debias by subtracting the estimated the NDE from the TE, retaining only the TIE. The final prediction is:
@@ 3.1 Experimental Setup
Dataset. We use two datasets: 1) ICBHI Respiratory Sound Database [22], for ID set, and 2) SPRSound [23], the Shanghai Jiao Tong University (SJTU) Pediatric Respiratory Sound Database, for OOD test set. First, ICBHI is annotated into four classes. Following the BTS [3], we binarize age into pediatric and adult groups and retain all other metadata as per the original ICBHI dataset specification. Second, SPRSound originally contains seven classes; to align with ICBHI, we merge crackle-related labels into crackle and combine Stridor and Rhonchi into wheeze. We also use the inter-patient-level validation set only for OOD test set. Comprehensive details are summarized in Table 1.
Training Details. Following BTS [3], we extract respiratory cycles, standardize them to 8 s, and resample the audio to 48 kHz. Text descriptions are capped at 64 tokens, which are sufficient to encode all text metadata without truncation. In line with BTS, we use all metadata and avoid relying solely on common attributes (e.g., sex) across datasets, ensuring the best ID performance.
For counterfactual debiasing, we set $\lambda_{\text{CE}}=1.0$ and $\lambda_{\text{KL}}=1.0$, respectively. We vary the value of $\alpha$ from 0 to 1 in increments of 0.1 to analyze the effect of $\alpha$ on overall performance. Counterfactual metadata augmentation is employed with a probability $p$ of 0.25. For adversarial debiasing, the adversary is applied with the coefficient of 1.0, while the location and device discriminator losses are weighted by $\lambda_{\text{location}}=0.01$ and $\lambda_{\text{device}}=0.1$, respectively. We fine-tune using the AdamW optimizer [29], with a learning rate of $5\times 10^{-5}$ for all parameters. It is further decayed via cosine scheduling over 30 epochs with a batch size of 8.
Evaluation Protocol. We use Specificity ($S_{p}$), Sensitivity ($S_{e}$), and their arithmetic mean (ICBHI Score) [22]. $S_{p}$ is the proportion of normal cases and $S_{e}$ is that of abnormal cases correctly classified. Results are reported as the mean and variance over five runs. We use PyTorch [30] and a single NVIDIA RTX 3090 for all experiments.
Table 1: Details of the ICBHI and SPRSound datasets. L/R denotes left or right. ‘Both’ label is co-occurrence of crackles and wheezes.
@@
Class Dist. Label / Ratio
Train Valid Overall
2063 (49.81%) 1579 (57.29%) 3642 (52.80%)
1215 (29.33%) 649 (23.55%) 1864 (27.02%)
501 (12.10%) 385 (13.97%) 886 (12.84%)
363 (8.76%) 143 (5.19%) 506 (7.34%)
@@
Class Dist. Label / Ratio
Train Valid Overall
5159 (78.14%) 1040 (72.78%) 6199 (77.19%)
961 (14.56%) 83 (5.81%) 1044 (13.00%)
452 (6.85%) 305 (21.34%) 757 (9.43%)
30 (0.45%) 1 (0.70%) 31 (0.39%)
@@ 3.2.1 Main Results
Table 2: Main Results on ID (ICBHI) and OOD (SPRSound) settings. Best results are in boldface and the second-best results are underlined.
Criteria Method Venue ID (In-Distribution) OOD (Out-of-Distribution)
@@
As shown in Table 2, our framework consistently enhances performance in both ID and OOD settings. It shows improved robustness to distribution shifts while preserving strong accuracy within the source domain. By contrast, conventional multimodal approaches such as BTS, which fail to account for spurious correlations, perform well under in-distribution but deteriorate markedly under distribution shifts. Unimodal baselines perform even worse, highlighting the efficacy of our debiasing strategy. Overall, the proposed elements contribute to performance improvement across heterogeneous data distributions compared to existing methodologies.
@@ 3.2.2 Ablation Studies
Table 3: Ablation studies on ID (ICBHI) and OOD (SPRSound) settings. Components are denoted as: (a) counterfactual reasoning, (b) adversarial debiasing, and (c) counterfactual metadata augmentation. For (b), (c), and Full, the best results across $\alpha$ are reported.
Components ID (In-Distribution) OOD (Out-of-Distribution)
@@ 3.2.3 Comparative Studies
Table 4: Comparative studies of debiasing attributes on ID (ICBHI) and OOD (SPRSound) settings.
Combinations ID (In-Distribution) OOD (Out-of-Distribution)
@@
Next, we investigate the effect of debiasing metadata attributes. Debiasing age leads to moderate improvements for evaluations conducted on both ID and OOD scenarios, with consistently balanced scores. Debiasing sex improves ID stability but considerably degrades OOD performance, suggesting that although this attribute is correlated with label bias, it still conveys information that generalizes across hospitals. Debiasing recording location and device yields relatively balanced outcomes, with strong ID performance while preserving OOD robustness. Further, debiasing age and sex does not generalize well OOD, and debiasing three components of age, location, and device significantly weakens OOD results. Therefore, these findings suggest that our strategy, debiasing both device and location simultaneously, provides the most reliable predictions.
@@ 3.2.4 Parameter Analysis
Fig. 3: Parameter analysis across different values of $\alpha$ on ID (ICBHI) and OOD (SPRSound) settings.
We analyze the effect of the debiasing coefficient $\alpha$, which controls the degree to which the NDE is subtracted from the TE during inference. As shown in Figure 3, while ID performance remains relatively stable across different $\alpha$ values, we observe a deterioration in OOD performance as $\alpha$ increases. This decline arises because excessive suppression also removes informative signals for generalization.
Interestingly, the strongest generalization is achieved at $\alpha=0$, where inference depends solely on the TE. This outcome suggests that our framework already debiases the TE effectively during training, obviating the need for explicit NDE subtraction at test time. In practice, setting $\alpha=0$ further reduces inference costs and eliminates the burden of tuning an additional hyperparameter, making it a cost-efficient and clinically pragmatic choice.

### v2

@@ Abstract
Multimodal respiratory sound classification offers promise for early pulmonary disease detection by integrating bioacoustic signals with patient metadata. Nevertheless, current approaches remain vulnerable to spurious correlations from attributes such as age, sex, or acquisition device, which hinder their generalization, especially under distribution shifts across clinical sites. To this end, we propose a counterfactual adversarial debiasing framework. First, we employ a causal graph-based counterfactual debiasing methodology to suppress non-causal dependencies from patient metadata. Second, we introduce adversarial debiasing to learn metadata-insensitive representations and reduce metadata-specific biases. Third, we design counterfactual metadata augmentation to mitigate spurious correlations further and strengthen metadata-invariant representations. By doing so, our method consistently outperforms strong baselines in evaluations under both in-distribution and distribution shifts. Code is available at https://github.com/RSC-Toolkit/BTS-CARD.
@@ 1 Introduction
Respiratory sound classification (RSC) provides a non-invasive and cost-effective method for the early detection of pulmonary diseases, offering a practical alternative to manual auscultation [21, 1, 4]. Clinicians typically base their diagnostic decisions on heterogeneous information sources, including patient metadata [4, 9, 11]. Motivated by this clinical reasoning process, recent advances in RSC have leveraged multimodal learning to integrate respiratory audio with auxiliary metadata (age, sex, stethoscope, etc.) [4].
@@
Causal inference provides a principled framework to address such limitations by estimating the effect of specific variables while controlling for confounders [18, 22]. This perspective is especially valuable in multimodal learning, where multimodal integration can make it more challenging to identify the true causal mechanisms, while also amplifying modality-driven biases [2, 16]. In RSC, metadata often exhibit skewed distributions across diagnostic categories, leading models to rely on superficial correlations instead of clinically meaningful features [24, 27]. Counterfactual debiasing mitigates this by disentangling causal signals from spurious ones, thereby enhancing in-distribution (IND) performance and out-of-distribution (OOD) robustness [22, 3].
Generalization is critical for clinical AI systems to ensure reliable performance across heterogeneous hospitals and patient populations [27, 8, 7]. Beyond causal methods, robustness under OOD conditions requires learning representations that remain stable across variations in demographics, acquisition devices, and clinical protocols [23]. Adversarial debiasing, originally introduced for algorithmic fairness, can be adapted to suppress hospital-specific attributes by jointly optimizing the main prediction task and adversarial domain classifiers [29, 13, 28]. Such training encourages metadata-agnostic representations and improves cross-site generalization [13]. However, existing research on RSC has overlooked the integration of counterfactual reasoning with adversarial debiasing to improve robustness under distribution shifts.
Therefore, we propose BTS-CARD, a novel counterfactual adversarial debiasing framework for multimodal RSC (see Figure 1). First, we introduce a counterfactual debiasing that suppresses non-causal associations, yielding more reliable predictions in both IND (ICBHI [19]) and OOD (SPRSound [30]) settings. Second, we enhance generalization via adversarial debiasing, which promotes metadata-invariant representations and prevents encoding acquisition device and recording location. Third, we devise counterfactual metadata augmentation to further mitigate spurious correlations from patient metadata and strengthen metadata-agnostic representations by neutralizing sensitive attributes. Extensive experiments demonstrate that our method consistently outperforms strong baselines, validating its robustness and transferability.
Fig. 1: An Overview of Our Counterfactual Adversarial Debiasing Framework for Multimodal RSC, BTS-CARD.
@@ 2.2 Task Formulation
Debiased Multimodal RSC. From a causal perspective, we decompose the influence of metadata into two pathways: the spurious direct path $t\rightarrow Y_{t,m}$ and the informative indirect path $(a,t)\rightarrow m\rightarrow Y_{t,m}$. Our framework employs counterfactual debiasing, illustrated by contrasting the top (factual) and bottom (counterfactual) graphs in the Counterfactual Inference of Figure 1, to disentangle causal effects from confounding. Combined with adversarial debiasing and counterfactual metadata augmentation, our approach further learns representations resilient to metadata variance, thereby improving accuracy and robustness against distribution shifts.
@@ 2.3 Counterfactual Adversarial Debiasing Framework
In this section, we present the counterfactual adversarial framework for multimodal RSC that builds upon the BTS model [4].
@@
where $Y_{t,m}$ denotes the prediction under text derived by metadata prompt $t$ [4] and multimodal representation $m$, while $Y_{t^{*},m^{*}}$ corresponds to their counterfactual counterparts. Here, multimodal counterfactual features are instantiated using a dummy vector $m^{*}$, fixed to a constant value of 1 based upon empirical validation. We use RUBi [2] for fusion, given its established effectiveness in causal modeling. Next, we estimate the NDE of metadata text using Clinical TinyBERT [20], which is not only efficient, but also easily adaptable to new tasks via fine-tuning. The NDE isolates the spurious shortcut $T\to Y$, and the TIE is then obtained by subtracting NDE from TE:
@@
Adversarial Debiasing. To promote metadata-agnostic representations for robustness across different clinical environments, we apply adversarial debiasing [29] to the NDE. To mitigate deployment bias, both recording devices and locations are targeted as they vary across clinical sites.
Debiasing is restricted to the natural direct effect (NDE), with logits $Y_{t}$ fed to an adversary $U$ to yield representation $z$. A Gradient Reversal Layer (GRL, $\mathcal{R}$) between $z$ and $D_{a}$ reverses gradients to promote metadata-insensitive features, with both tasks optimized via CE. Thus, with $D_{a}$, the adversarial discriminator for the metadata attribute $a\in\{\text{location},\text{device}\}$ and $l_{a}$, its corresponding ground-truth label, the objective is as follows:
@@
$\text{where}\>L_{\text{CE}_{\text{NDE}}}$ $=\text{CE}(Y_{t},y)\>\text{and}\>L_{a}=\text{CE}(D_{a}(\mathcal{R}(z)),l_{a}).$
@@
From a causal perspective, this intervention suppresses the spurious shortcut $T\to Y$ while leaving the TE model intact to preserve both direct and indirect effects [22]. This methodology enforces insensitivity to non-causal attributes without discarding meaningful pathways. It parallels the logic of $do$-calculus [18], where interventions isolate causal effects by severing spurious dependencies. Replacing sensitive metadata in this manner thus mimics a $do$-operation that removes the $T\to Y$ link while preserving potential mediating paths to better reflect the hypothesized causal structure.
@@
Thus, with $\lambda_{\text{CE}}$ and $\lambda_{\text{KL}}$ controlling the weights of CE and KL regularization, the overall training loss is:
@@
At inference, we debias by subtracting the estimated NDE from the TE, retaining only the TIE. The final prediction is:
@@ 3.1 Experimental Setup
Dataset. We use two datasets: 1) ICBHI Respiratory Sound Database [19], for IND set, and 2) SPRSound [30], the Shanghai Jiao Tong University (SJTU) Pediatric Respiratory Sound Database, for OOD test set. First, ICBHI is annotated into four classes. Following the BTS [4], we binarize age into pediatric and adult groups and retain all other metadata as per the original ICBHI dataset specification. Second, SPRSound originally contains seven classes; to align with ICBHI, we merge crackle-related labels into crackle and combine Stridor and Rhonchi into wheeze. We also use the inter-patient-level validation set only for OOD test set. Comprehensive details are summarized in Table 1.
Training Details. Following BTS [4], we extract respiratory cycles, standardize them to 8 s, and resample the audio to 48 kHz. Text descriptions are capped at 64 tokens, which are sufficient to encode all text metadata without truncation. In line with BTS, we use all metadata and avoid relying solely on common attributes (e.g., sex) across datasets, ensuring the best IND performance.
For counterfactual debiasing, we set $\lambda_{\text{CE}}=1.0$ and $\lambda_{\text{KL}}=1.0$, respectively. We vary the value of $\alpha$ from 0 to 1 in increments of 0.1 to analyze the effect of $\alpha$ on overall performance. During training, counterfactual metadata augmentation is employed by independently replacing each metadata attribute with a neutral placeholder with a probability $p$ of 0.25. For adversarial debiasing, the adversary is applied with the coefficient of 1.0, while the location and device discriminator losses are weighted by $\lambda_{\text{location}}=0.01$ and $\lambda_{\text{device}}=0.1$, respectively. We fine-tune using the AdamW optimizer [14], with a learning rate of $5\times 10^{-5}$ for all parameters. It is further decayed via cosine scheduling over 30 epochs with a batch size of 8.
Evaluation Protocol. We use Specificity ($S_{p}$), Sensitivity ($S_{e}$), and their arithmetic mean (ICBHI Score) [19]. $S_{p}$ is the proportion of normal cases, and $S_{e}$ is that of abnormal cases correctly classified. Results are reported as the mean and variance over five runs. We use PyTorch [17] and a single NVIDIA RTX 3090 for all experiments.
Table 1: Details of the ICBHI and SPRSound datasets. L/R denotes left or right. ‘Both’ label is co-occurrence of crackle and wheeze.
@@
Class Dist. Label / Ratio Train Valid Overall 2063 (49.81%) 1579 (57.29%) 3642 (52.80%) 1215 (29.33%) 649 (23.55%) 1864 (27.02%) 501 (12.10%) 385 (13.97%) 886 (12.84%) 363 (8.76%) 143 (5.19%) 506 (7.34%)
@@
Class Dist. Label / Ratio Train Valid Overall 5159 (78.14%) 1040 (72.78%) 6199 (77.19%) 961 (14.56%) 83 (5.81%) 1044 (13.00%) 452 (6.85%) 305 (21.34%) 757 (9.43%) 30 (0.45%) 1 (0.70%) 31 (0.39%)
@@ 3.2.1 Main Results
Table 2: Main Results on IND (ICBHI) and OOD (SPRSound) settings. Best results are in boldface and the second-best results are underlined.
Criteria Method Venue IND (In-Distribution) OOD (Out-of-Distribution)
@@
As shown in Table 2, our framework consistently enhances performance in both IND and OOD settings. By contrast, conventional multimodal approaches such as BTS, which fail to account for spurious correlations, perform well under in-distribution but deteriorate markedly under distribution shifts. Unimodal baselines perform even worse, highlighting the efficacy of our debiasing methodology. Overall, the proposed elements contribute to performance improvement across heterogeneous data distributions compared to existing methodologies.
@@ 3.2.2 Ablation Studies
Table 3: Ablation studies on IND (ICBHI) and OOD (SPRSound) settings. Components are denoted as: (a) counterfactual debiasing, (b) adversarial debiasing, and (c) counterfactual metadata augmentation. For (b), (c), and Full, the best results across $\alpha$ are reported.
Components IND (In-Distribution) OOD (Out-of-Distribution)
@@ 3.2.3 Comparative Studies
Table 4: Comparative studies of debiasing attributes on IND (ICBHI) and OOD (SPRSound) settings.
Combinations IND (In-Distribution) OOD (Out-of-Distribution)
@@
Next, we investigate the effect of debiasing metadata attributes. Debiasing age leads to moderate improvements for evaluations conducted on both IND and OOD scenarios, with consistently balanced scores. Debiasing sex improves IND stability but considerably degrades OOD performance, suggesting that although this attribute is correlated with label bias, it still conveys information that generalizes across hospitals. Debiasing recording location and device yields relatively balanced outcomes, with strong IND performance while preserving OOD robustness. Further, debiasing age and sex does not generalize well OOD, and debiasing three components of age, location, and device significantly weakens OOD results. Therefore, these findings suggest that our methodology, debiasing both device and location simultaneously, provides the most reliable predictions.
@@ 3.2.4 Parameter Analysis
Fig. 3: Parameter analysis across different values of $\alpha$ on IND (ICBHI) and OOD (SPRSound) settings.
We analyze the effect of the debiasing coefficient $\alpha$, which controls the degree to which the NDE is subtracted from the TE during inference. As shown in Figure 3, IND performance remains relatively stable across $\alpha$, whereas OOD performance deteriorates as $\alpha$ increases. This trade-off indicates that excessive suppression removes informative signals for generalization, prioritizing bias removal over robustness and resulting in saturated IND performance at the cost of degraded OOD generalization.
Interestingly, the strongest generalization is achieved at $\alpha=0$, where inference relies solely on the TE. This suggests that our framework effectively debiases the TE during training, alleviating the need for an explicit trade-off between IND accuracy and OOD robustness at test time. In practice, setting $\alpha=0$ reduces inference costs and eliminates the need to tune an additional hyperparameter, making it a cost-efficient and clinically pragmatic choice.